A trolley, while going down an inclined plane, has an acceleration of 2 cm s-2. What will be its velocity 3 s after the start?

Grade:12th pass

1 Answers

Harshit Singh
2 years ago
Dear Student

Initial Velocity of the trolley, u= 0 cms-1
Acceleration, a= 2 cms-2
Time, t= 3 s
It is known that final velocity, v= u + at = 0 + 2 x 3 cms-1
Therefore, the velocity of train after 3 seconds is 6 cms-1

Thanks